# Part-aware Shape Generation with Latent 3D Diffusion of Neural Voxel Fields

### v3

@@ Abstract
This paper introduces a novel latent 3D diffusion model for generating neural voxel fields with precise part-aware structures and high-quality textures. In comparison to existing methods, this approach incorporates two key designs to guarantee high-quality and accurate part-aware generation. On one hand, we introduce a latent 3D diffusion process for neural voxel fields, allowing generation at significantly higher resolutions to capture rich textural and geometric details accurately. On the other hand, a part-aware shape decoder is introduced to integrate the part codes into the neural voxel fields, guiding accurate part decomposition and producing high-quality rendering results. Importantly, part-aware learning establishes structural relationships to generate texture information for similar regions, thereby facilitating high-quality rendering results. We evaluate our approach across eight different data classes through extensive experimentation and comparisons with state-of-the-art methods. The results demonstrate that our proposed method has superior generative capabilities in part-aware shape generation, outperforming existing state-of-the-art methods. Moreover, we have conducted image- and text-guided shape generation via the conditioned diffusion process, showcasing the advanced potential in multi-modal guided shape generation.
@@ I Introduction
We propose a latent 3D diffusion model to generate neural voxel fields with accurate structural details in this paper. As shown in Fig 1, our network is composed of two parts, a latent 3D diffusion model of neural voxel fields and a part-aware shape decoder. Specifically, the latent 3D diffusion model utilizes the advantage of 3D convolution and employs the 3D denoising process in a latent space with a much lower resolution. On the other hand, the part-aware shape decoder incorporates the part code with the neural voxel field to learn part-aware relationships, guiding the precise generation of part-aware shapes and synthesizing high-quality rendering results.
Benefiting from the latent diffusion design, our model allows for shape generation with a much higher resolution than previous work (e.g., Matthias [18] and [19]). In particular, our method achieves $96^{3}$ resolution which is sufficient for representing rich geometric details of most common-seen objects. Furthermore, we integrate cross- and self-attention mechanisms into the part-aware decoder to establish robust dependencies between the part code and the voxel field. This integration aids in modeling the inherent relationships among various parts, significantly enhancing the precision of part segmentation. More importantly, learning part-aware information enhance the correlations between different regions, facilitating the geometry reconstruction and texture modeling in similar parts.
@@
Figure 2: The proposed method comprises two main modules: a latent 3D diffusion model and a part-aware shape decoder. The latent 3D diffusion model facilitates high-resolution generation of neural voxel fields, while the part-aware shape decoder enables the generation of parts-aware results for rendering. The two modules are trained in an end-to-end fashion.
@@
The diffusion model admits multi-modal prompt for shape generation. For example, we have implemented image-conditioned diffusion enabling single-image guided generation. Concretely, a convolutional encoder is employed to extract features from the reference image. Subsequently, these obtained features are combined with the multi-scale features of the denoising UNet. As a result, the image information can effectively guide the shape generation. Moreover, we perform text-conditioned diffusion for text-guided shape generation. We utilize the general language model CLIP [20] to extract embedding for the text input, and perform cross-attention between the text embedding and the multi-scale features of the denoising UNet.
@@ II Related Work
Diffusion Models for Shape Generation. Recent developments in diffusion models [21] have enabled the synthesis of 3D shapes through a variety of representations, including point cloud [22, 23], SDF [24], mesh [1], and neural field [12, 25, 26, 27, 28]. [29] proposes to utilize diffusion models to generate a compact wavelet representation with a pair of coarse and detail coefficient volumes. 3DShape2VecSet [30] introduces a novel representation for neural fields designed for generative diffusion models, which can encode 3D meshes or point clouds as neural fields. DiffFacto [31] learns the distribution of point clouds with part-level control including part-level sampling, mixing and interpolation. EXIM [32] use the text prompt as guidance, generating corresponding truncated signed distance field with diffusion models. However, most of above methods must adopt 3D input data as supervision, which is not available easily. Differently, this paper focus on generating neural voxel fields with differentiable rendering, enabling multi-view image as supervision.
@@ III Proposed Method
In this section, we introduce a latent 3D diffusion model to generate high-resolution part-aware neural voxel fields. As illustrated in Fig. 2, our approach consists of a latent 3D diffusion model and a part-aware shape decoder. Specifically, the latent 3D diffusion model compresses the object fields into the latent space with a much smaller spatial resolution, and then adopts the 3D UNet to conduct the diffusion process. This implementation facilitates the high resolution of the voxel field, allowing the high-quality rendering results. On the other hand, the part-aware shape decoder integrates the part codes into the voxel field, establishing relationships between different parts of the voxel field, and produces RGB images and semantic part maps through ray casting.
@@ III-A1 Neural Voxel Fields
Neural radiance fields (NeRFs) [47] represent 3D object implicitly and map the 3D position $\boldsymbol{x}$ and view direction $\boldsymbol{d}$ to the density $\sigma$ and color emission $\boldsymbol{c}$, obtaining the novel view image via volume rendering. Based on NeRFs, DVGO [48] learns neural voxel fields to represent the object’s density and color. Specifically, there are a density voxel field and a feature voxel field, which are represented by $\mathbf{V}^{\sigma}$ and $\mathbf{V}^{\boldsymbol{c}}$. We can acquire the density and color feature of position $\boldsymbol{x}$ by querying from $\mathbf{V}^{\sigma}$ and $\mathbf{V}^{\boldsymbol{c}}$ directly, and the color feature and view direction are feed into multilayer perceptron (MLP) layers to get the color emission $\boldsymbol{c}$.
@@ III-B Latent 3D Diffusion Model
As shown in Fig. 2, the proposed latent 3D diffusion model comprises a 3D autoencoder and a 3D denoising UNet. The 3D autoencoder has an encoder for compressing the voxel field $\mathbf{V}^{\boldsymbol{f}}_{0}$ to the latent code $\mathbf{L}_{0}$ and a decoder for recovering the original voxel field from $\mathbf{L}_{0}$, i.e. $\mathbf{L}_{0}=\mathcal{E}(\mathbf{V}^{\boldsymbol{f}}_{0}),\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}=\mathcal{D}(\mathbf{L}_{0})$. Here, $\mathcal{E}$ and $\mathcal{D}$ represent the encoder and decoder of the 3D autoencoder; $\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}$ is the reconstructed voxel field. In practice, the resolution of $\mathbf{L}_{0}$ is 64$\times$ lower than $\mathbf{V}^{\boldsymbol{f}}_{0}$. After obtaining the latent code, we adopt a 3D UNet architecture to conduct the diffusion process in the latent space. Similar to Eq. 3, we sample the noisy code $\mathbf{L}_{t}$ by adding the normal noise $\boldsymbol{\epsilon}$ to $\mathbf{L}_{0}$. Then, the noisy code is fed into the 3D UNet with the time $t$ together, outputting the estimations of $\mathbf{L}_{0}$ and $\boldsymbol{\epsilon}$. This process is formulated by:
@@
$\mathbf{L}_{\boldsymbol{\phi}},\boldsymbol{\epsilon}_{\boldsymbol{\phi}}=\text{UNet}_{\boldsymbol{\phi}}(\mathbf{L}_{t},t),$
@@
For conditioned generation, we utilize the classifier-free [52] manner to synthesize shapes corresponding to the conditioned input. Specifically, the conditioned input is fed into a feature extractor, resulting in a feature embedding. We adopt the feature embedding as an additional input of the 3D UNet, thus we obtain the $\mathbf{L}_{\boldsymbol{\phi}}$ and $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}$ by: $\mathbf{L}_{\boldsymbol{\phi}},\boldsymbol{\epsilon}_{\boldsymbol{\phi}}=\text{UNet}_{\boldsymbol{\phi}}(\mathbf{L}_{t},t,\mathbf{e})$, where $\mathbf{e}$ is the feature embedding of the conditioned input.
3D UNet vs. 2D UNet. Actually, though the latent code $\mathbf{L}_{t}$ is a 4D tensor, it is possible to apply the 2D convolutional operator to this type of data. In practice, we can reshape $\mathbf{L}_{t}$ into a 3D tensor, i.e. $\text{reshape}(\mathbb{R}^{C\times X\times Y\times Z})\rightarrow\mathbb{R}^{CX\times Y\times Z}$, thus adopting 2D UNet to model the denoising process. However, we found that the loss curve did not converge and it failed to obtain satisfactory results (details can be found in Sec. IV-C). We assume that the 2D UNet architecture hardly learns the 3D consistency of the voxel field, resulting in a perturbed result. In contrast, the 3D UNet architecture provides reasonable performance, demonstrating the necessity of adopting 3D UNet for denoising the neural voxel field.
@@ III-C Part-aware Shape Decoder
After the denoising process of the latent 3D diffusion model, we can have a neural voxel field containing the density and color features. However, it can not provide the structural information and rendering results. Therefore, we propose a part-aware shape decoder to generate part-aware details and rendering results.
As shown in Fig. 2, to model the part-aware information, the shape decoder integrates the part code into the voxel field and utilizes the part attention to build the relations between different parts of the voxel field. Specifically, the part code is a learnable embedding and can be represented by $\boldsymbol{z}\in\mathbb{R}^{K\times D}$, where $K$ is the number of parts, $D$ is the dimension of the embedding. To incorporate the part code into the neural field, we first project the reconstructed voxel field $\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}$ to have the same dimension as $\boldsymbol{z}$, and then flatten $\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}$ to be a 2D tensor like $\mathbb{R}^{XYZ\times D}$. Thus, we can utilize the multi-head cross-attention [53] to build a strong dependency between the part code and the voxel field. Moreover, to divide the voxel field accurately, we employ a self-attention to model the long-range relations between the different parts of the voxel field. After the cross-attention and self-attention, we reshape the tensor back to be a 4D voxel field. We formulate this process by the following equation:
@@ III-D End-to-end Training with Gradient Skip
$\nabla_{\boldsymbol{\phi}}{\mathcal{L}_{rend}}=\frac{\partial\mathcal{L}_{rend}}{\partial\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}}\frac{\partial\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}}{\partial\mathbf{L}_{\boldsymbol{\phi}}}\frac{\mathbf{L}_{\boldsymbol{\phi}}}{\partial\boldsymbol{\phi}},$ (9)
@@
$\nabla_{\boldsymbol{\phi}}{\mathcal{L}_{rend}}=\frac{\partial\mathcal{L}_{rend}}{\partial\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}}\frac{\mathbf{L}_{\boldsymbol{\phi}}}{\partial\boldsymbol{\phi}}.$ (10)
@@
Except for the main training objective, we adopt two additional regularization loss functions. First, we conduct the total variation on the neural voxel fields to restrain the noise generating:
@@ IV-A1 Dataset
We collect eight categories of 3D objects from ShapeNet [56], including Chair, Table, Airplane, Car, Lamp, Guitar, Laptop, and Pistol. For each category, we random sample 500 objects, with a split of 450 for training and 50 for testing purposes. Utilizing BlenderProc [57], we render 100 views of each object at a resolution of 800$\times$800 resolution. With the posed images, we use the public code base [48] to generate neural voxel fields. For text-guided shape generation, we get the text labels of Chair and Table from [58].
@@ IV-A2 Implementation Details
The resolution of the voxel field is 963, and the resolution of the latent code is 64$\times$ lower than the voxel field. We train the model for 100,000 iterations and we do not train the part-aware shape decoder for the first 20,000 iterations to warm up the latent 3D diffusion model. We set $K=4$ and $D=128$ for the part code. For each iteration, we sample 8192 rays of 5 random views to calculate the rendering loss. For the architecture of 3D UNet, we set the base channel to 96 and the channel multiplier is [1, 2, 4], resulting in 144M parameters of the whole model. We train the model in a single RTX 3090 GPU with batch size 2, taking up around 15, [^0] M GPU memory cost.

### v5

@@ Abstract
This paper introduces a novel latent 3D diffusion model for generating neural voxel fields with precise part-aware structures and high-quality textures. In comparison to existing methods, this approach incorporates two key designs to guarantee high-quality and accurate part-aware generation. On one hand, we introduce a latent 3D diffusion process for neural voxel fields, incorporating part-aware information into the diffusion process and allowing generation at significantly higher resolutions to capture rich textural and geometric details accurately. On the other hand, a part-aware shape decoder is introduced to integrate the part codes into the neural voxel fields, guiding accurate part decomposition and producing high-quality rendering results. Importantly, part-aware learning establishes structural relationships to generate texture information for similar regions, thereby facilitating high-quality rendering results. We evaluate our approach across eight different data classes through extensive experimentation and comparisons with state-of-the-art methods. The results demonstrate that our proposed method has superior generative capabilities in part-aware shape generation, outperforming existing state-of-the-art methods. Moreover, we have conducted image- and text-guided shape generation via the conditioned diffusion process, showcasing the advanced potential in multi-modal guided shape generation.
@@ I Introduction
We propose a latent 3D diffusion model to generate neural voxel fields with accurate structural details in this paper. As shown in Fig 1, our network consists of two parts, a latent 3D diffusion model of neural voxel fields and a part-aware shape decoder. Specifically, the latent 3D diffusion model utilizes the advantage of 3D convolution and employs the 3D denoising process in a latent space with a much lower resolution. On the other hand, the part-aware shape decoder incorporates the part code with the neural voxel field to learn part-aware relationships, guiding the precise generation of part-aware shapes and synthesizing high-quality rendering results. Moreover, the part code is also incorporated into the diffusion model, which allows for learning the part-aware information in field generation.
Benefiting from the latent diffusion design, our model allows for shape generation with a much higher resolution than previous work (e.g., Matthias [18] and [19]). In particular, our method achieves $96^{3}$ resolution which is sufficient for representing rich geometric details of most common-seen objects. Furthermore, we integrate cross- and self-attention mechanisms into the part-aware decoder to establish robust dependencies between the part code and the voxel field. This integration aids in modeling the inherent relationships among various parts, significantly enhancing the precision of part segmentation. More importantly, learning part-aware information enhances the correlations between different regions, facilitating geometry reconstruction and texture modeling in similar parts.
@@
Figure 2: The proposed method comprises two main modules: a latent 3D diffusion model and a part-aware shape decoder. The latent 3D diffusion model facilitates the high-resolution generation of neural voxel fields, while the part-aware shape decoder enables the generation of parts-aware results for rendering. The two modules are trained jointly.
@@
The diffusion model admits multi-modal prompts for shape generation. For example, we have implemented image-conditioned diffusion enabling single-image guided generation. Concretely, a convolutional encoder is employed to extract features from the reference image. Subsequently, these obtained features are combined with the multi-scale features of the denoising UNet. As a result, the image information can effectively guide the shape generation. Moreover, we perform text-conditioned diffusion for text-guided shape generation. We utilize the general language model CLIP [20] to extract embedding for the text input, and perform cross-attention between the text embedding and the multi-scale features of the denoising UNet.
@@ II Related Work
Diffusion Models for Shape Generation. Recent developments in diffusion models [21] have enabled the synthesis of 3D shapes through a variety of representations, including point cloud [22, 23, 24], SDF [25], mesh [1], and neural field [12, 26, 27, 28, 29]. [30] proposes to utilize diffusion models to generate a compact wavelet representation with a pair of coarse and detail coefficient volumes. 3DShape2VecSet [31] introduces a novel representation for neural fields designed for generative diffusion models, which can encode 3D meshes or point clouds as neural fields. DiffFacto [32] learns the distribution of point clouds with part-level control including part-level sampling, mixing and interpolation. EXIM [33] use the text prompt as guidance, generating corresponding truncated signed distance field with diffusion models. However, most of above methods must adopt 3D input data as supervision, which is not available easily. Differently, this paper focus on generating neural voxel fields with differentiable rendering, enabling multi-view image as supervision.
@@ III Proposed Method
In this section, we introduce a latent 3D diffusion model to generate high-resolution part-aware neural voxel fields. As illustrated in Fig. 2, our approach consists of a latent 3D diffusion model and a part-aware shape decoder. Specifically, the latent 3D diffusion model compresses the object fields into the latent space with a much smaller spatial resolution, and then adopts the 3D UNet to conduct the diffusion process. This implementation facilitates the high resolution of the voxel field, allowing the high-quality rendering results. Moreover, the part code is fed into the diffusion model to incorporate part-aware information into field generation. On the other hand, the part-aware shape decoder integrates the part codes into the voxel field, establishing relationships between different parts of the voxel field, and produces RGB images and semantic part maps through ray casting.
Figure 3: Detailed architecture of the 3D UNet. The 3D autoencoder does not have the modules in the dashed purple rectangular box.
@@ III-A1 Neural Voxel Fields
Neural radiance fields (NeRFs) [48] represent 3D objects implicitly and map the 3D position $\boldsymbol{x}$ and view direction $\boldsymbol{d}$ to the density $\sigma$ and color emission $\boldsymbol{c}$, obtaining the novel view image via volume rendering. Based on NeRFs, DVGO [49] learns neural voxel fields to represent the object’s density and color. Specifically, there are a density voxel field and a feature voxel field, which are represented by $\mathbf{V}^{\sigma}$ and $\mathbf{V}^{\boldsymbol{c}}$. We can acquire the density and color feature of position $\boldsymbol{x}$ by querying from $\mathbf{V}^{\sigma}$ and $\mathbf{V}^{\boldsymbol{c}}$ directly, and the color feature and view direction are feed into multilayer perceptron (MLP) layers to get the color emission $\boldsymbol{c}$.
@@ III-B Latent 3D Diffusion Model
As shown in Fig. 2, the proposed latent 3D diffusion model comprises a 3D autoencoder and a 3D denoising UNet. The 3D autoencoder has an encoder for compressing the voxel field $\mathbf{V}^{\boldsymbol{f}}_{0}$ to the latent code $\mathbf{L}_{0}$ and a decoder for recovering the original voxel field from $\mathbf{L}_{0}$, i.e. $\mathbf{L}_{0}=\mathcal{E}(\mathbf{V}^{\boldsymbol{f}}_{0}),\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}=\mathcal{D}(\mathbf{L}_{0})$. Here, $\mathcal{E}$ and $\mathcal{D}$ represent the encoder and decoder of the 3D autoencoder; $\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}$ is the reconstructed voxel field. In practice, the resolution of $\mathbf{L}_{0}$ is 64$\times$ lower than $\mathbf{V}^{\boldsymbol{f}}_{0}$. After obtaining the latent code, we adopt a 3D UNet architecture to conduct the diffusion process in the latent space. Our 3D UNet not only takes the noisy latent code and time as input, but also introduces the part code to incorporate structural information. In practice, the part code is a learnable embedding and can be represented by $\boldsymbol{z}\in\mathbb{R}^{K\times D}$, where $K$ is the number of parts, $D$ is the dimension of the embedding. Similar to Eq. 3, we sample the noisy code $\mathbf{L}_{t}$ by adding the normal noise $\boldsymbol{\epsilon}$ to $\mathbf{L}_{0}$. Then, the noisy code is fed into the 3D UNet with the time $t$ and part code $z$ together, outputting the estimations of $\mathbf{L}_{0}$ and $\boldsymbol{\epsilon}$. This process is formulated by:
@@
$\mathbf{L}_{\boldsymbol{\phi}},\boldsymbol{\epsilon}_{\boldsymbol{\phi}}=\text{UNet}_{\boldsymbol{\phi}}(\mathbf{L}_{t},t,z),$
@@
For the conditioned generation, we utilize the classifier-free [53] manner to synthesize shapes corresponding to the conditioned input. Specifically, the conditioned input is fed into a feature extractor, resulting in a feature embedding. We adopt the feature embedding as an additional input of the 3D UNet, thus we obtain the $\mathbf{L}_{\boldsymbol{\phi}}$ and $\boldsymbol{\epsilon}_{\boldsymbol{\phi}}$ by: $\mathbf{L}_{\boldsymbol{\phi}},\boldsymbol{\epsilon}_{\boldsymbol{\phi}}=\text{UNet}_{\boldsymbol{\phi}}(\mathbf{L}_{t},t,\mathbf{e})$, where $\mathbf{e}$ is the feature embedding of the conditioned input.
3D UNet vs. 2D UNet. Though the latent code $\mathbf{L}_{t}$ is a 4D tensor, it is possible to apply the 2D convolutional operator to this type of data. In practice, we can reshape $\mathbf{L}_{t}$ into a 3D tensor, i.e. $\text{reshape}(\mathbb{R}^{C\times X\times Y\times Z})\rightarrow\mathbb{R}^{CX\times Y\times Z}$, thus adopting 2D UNet to model the denoising process. However, we found that the loss curve did not converge and it failed to obtain satisfactory results (details can be found in Sec. IV-C). We assume that the 2D UNet architecture hardly learns the 3D consistency of the voxel field, resulting in a perturbed result. In contrast, the 3D UNet architecture provides reasonable performance, demonstrating the necessity of adopting 3D UNet for denoising the neural voxel field.
@@ III-C Part-aware Shape Decoder
After the denoising process of the latent 3D diffusion model, we can have a neural voxel field containing the density and color features. However, it can not provide the structural information and rendering results. Therefore, we propose a part-aware shape decoder to generate part-aware details and rendering results. As shown in Fig. 2, to model the part-aware information, the shape decoder integrates the part code $z$ into the voxel field and utilizes the attention mechanism to build the relations between different parts of the voxel field. To incorporate the part code into the neural field, we first project the reconstructed voxel field $\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}$ to have the same dimension as $\boldsymbol{z}$, and then flatten $\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}$ to be a 2D tensor like $\mathbb{R}^{XYZ\times D}$. Thus, we can utilize the multi-head cross-attention [54] to build a strong dependency between the part code and the voxel field. Moreover, to divide the voxel field accurately, we employ a self-attention to model the long-range relations between the different parts of the voxel field. After the cross-attention and self-attention, we reshape the tensor back to be a 4D voxel field. We formulate this process by the following equation:
@@ III-D End-to-end Training with Gradient Skip
$\nabla_{\boldsymbol{\phi}}{\mathcal{L}_{rend}}=\frac{\partial\mathcal{L}_{rend}}{\partial\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}}\frac{\partial\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}}{\partial\mathbf{L}_{\boldsymbol{\phi}}}\frac{\partial\mathbf{L}_{\boldsymbol{\phi}}}{\partial\boldsymbol{\phi}},$ (9)
@@
$\nabla_{\boldsymbol{\phi}}{\mathcal{L}_{rend}}=\frac{\partial\mathcal{L}_{rend}}{\partial\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}}\frac{\partial\mathbf{L}_{\boldsymbol{\phi}}}{\partial\boldsymbol{\phi}}.$ (10)
Since the shape of $\mathbf{L}_{\boldsymbol{\phi}}$ is different from $\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}$, we use trilinear interpolation to align their shapes for calculating the above equation.
@@
Except for the main training objective, we adopt two additional regularization loss functions. First, we conduct the total variation on the neural voxel fields to restrain the noise generation:
@@
III-E Detailed Architectures
We adopt the 3D UNet architecture following [57] and add the additional time and part code input branches. More specifically, the time and part code are projected to be a scale weight and a bias weight, and the two weights are used to transform the intermediate variable of the 3D convolution layer. The 3D autoencoder is also a 3D UNet architecture without additional branches. The detailed architecture of the 3D UNet is depicted in Fig. 3.
@@ IV-A1 Dataset
We collect eight categories of 3D objects from ShapeNet [58], including Chair, Table, Airplane, Car, Lamp, Guitar, Laptop, and Pistol. For each category, we random sample 500 objects, with a split of 450 for training and 50 for testing purposes. Utilizing BlenderProc [59], we render 100 views of each object at a resolution of 800$\times$800 resolution. With the posed images, we use the public code base [49] to generate neural voxel fields. For text-guided shape generation, we get the text labels of Chair and Table from [60]. For ground truth part labels, we use the shape segmentation method [61] to obtain the labeled shapes, and then use BlenderProc to render 2D segmentation maps.
@@ IV-A2 Implementation Details
The resolution of the voxel field is 963, and the resolution of the latent code is 64$\times$ lower than the voxel field. We train the model for 100,000 iterations and we do not train the part-aware shape decoder for the first 20,000 iterations to warm up the latent 3D diffusion model. We set $K=4$ and $D=128$ for the part code. For each iteration, we sample 8192 rays of 5 random views to calculate the rendering loss. For the architecture of 3D UNet, we set the base channel to 96 and the channel multiplier is [1, 2, 4], resulting in 144M parameters of the whole model. We train the model in a single RTX 3090 GPU with batch size 2, taking up around 15, [^0] M GPU memory cost. The 3D autoencoder is pretrained on the same NeRF dataset as the whole model, and we apply a Kullback-Leibler Divergence loss on the latent codes to regularize its distribution close to the normal distribution following [62].